# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v2

@@ 5.2 Gradient domination with $\zeta=2$
For a fixed $\epsilon\in\left(0,\frac{\lambda_{0}}{\lambda_{1}}\right)$, if $f$ satisfies Assn. 1, 2 and 3 and Assn. 4 with $\zeta=2$, $f^{*}=0$ and if $\mu:=\min_{t\in[T]}\mu(\theta_{t})>0$, then, GD-LS with $\eta_{\max}=\infty$, requires
$T\geq\frac{2}{\mu}\,\left[\lambda_{1}{f}(\theta_{0})+\lambda_{0}\,\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right]$
iterations to ensure that $f(\theta_{T})\leq\epsilon$.
@@ 6 SGD with Stochastic Line-search
When minimizing uniformly-smooth convex functions in the interpolation setting, Vaswani et al. (2019a) proved that SGD-SLS converges to the optimum at an $O(1/\epsilon)$ rate, matching GD and is faster than the standard $O(1/\epsilon^{2})$ rate for SGD (Bottou et al., 2018). Moreover, SGD-SLS does not require the knowledge of the global smoothness constant, making it an attractive choice in practice. Motivated by this and the results in Sec. 4, we analyze SGD-SLS for convex functions that satisfy Assn. 1, 2 and 3.
Since SLS involves an Armijo line-search for one (randomly chosen) function in each iteration, we can follow the same argument as in Lemma 1 and show that the step-size in each iteration is lower-bounded i.e. ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}\,+\lambda_{1}\,{f_{t}}(\theta_{t})}\right\}$. Given this result, we first prove the following lemma in App. E.
Lemma 3.
For a fixed $\epsilon>0$ and comparator $u$, assuming $f(\theta):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(\theta)$ where each $f_{i}$ is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0$, if $\Delta_{t}:=\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]$, $f(u)\leq f(\theta_{t})$ for all $t\in[T]$, and $C:=\frac{(2c-1)}{c\,\lambda_{1}}$, then, SGD-SLS guarantees that:
$\Delta_{T}$ $\leq\Delta_{0}-\sum_{t=0}^{T-1}\mathbb{E}\left[\min\left\{\eta_{\max},\frac{C}{f(\theta_{t})}\right\}\,[f(\theta_{t})-f(u)]\right]$
$+2\eta_{\max}\,\chi^{2}(u)\,T\,,$
where the expectation is over the randomness in all iterations $t\in[T]$ and $\chi^{2}(u):=\mathbb{E}_{i}[f_{i}(u)-\min f_{i}(\theta)]$ is a measure of the stochasticity at $u$.
Since logistic regression on linearly separable data is convex, satisfies Assn. 1, 2 and 3 and the interpolation condition, we follow a similar strategy as in Cor. 2 and prove the following corollary.
Corollary 5.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{1}{8}\right)$, SGD-SLS with $\eta_{\max}=O\left(\frac{(1-c)\,(2c-1)}{c}\,\frac{1}{\epsilon}\right)$ requires
$T\geq O\left(\frac{c}{(1-c)\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right)$
iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq 4\epsilon$.
Hence, SGD-SLS requires $O\left(\ln\left(\frac{1}{\epsilon}\right)^{2}\right)$ iterations (and hence gradient evaluations) to ensure convergence to an $\epsilon$ neighbourhood. For logistic regression on separable data, this is an exponential improvement over the $O(1/\epsilon)$ rate for SGD-SLS in Vaswani et al. (2019b) and SGD with a Polyak step-size in Loizou et al. (2021). Compared to the deterministic result in Sec. 4 which holds for the last-iterate, the above result only holds for the best-iterate. Furthermore, while Cor. 2 required $\eta_{\max}=\infty$, the above result requires $\eta_{\max}=O(\nicefrac{{1}}{{\epsilon}})$ in order to handle the stochasticity.
From a practical perspective, Vaswani et al. (2019b); Galli et al. (2024) have empirically demonstrated the effectiveness of SGD-SLS and its non-monotone variant for convex losses corresponding to linear and logistic regression, and non-convex losses with deep neural networks. In particular, for logistic regression on the mushrooms dataset where interpolation is satisfied (Vaswani et al., 2019b, Figure 3)Cor. 5 explains the linear convergence of SGD-SLS.
@@ C.1 Proofs for the Polyak Step-size
Following a proof similar to that for Thm. 2, for an arbitrary comparator $u$ s.t. $f(u)\leq(\theta_{T})$,
@@
We use the above result and prove the following corollary for logistic regression on separable data.
Corollary 6.
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, a fixed $\epsilon\in(0,f(\theta_{0}))$, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-f(\beta u^{*})}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
$T\geq\frac{4\,c^{2}}{(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
@@
For a scalar $\beta>0$,
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{\lambda_{\max}[X^{T}X]}{4n}\leq 1$. Moreover, $\nu=1$. We set $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$ implies that $f(\beta u^{*})\leq\epsilon$. For the $T$ defined in the theorem statement, consider two cases:
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, we can use the result in Eq. 23 in Thm. 4 with the comparator $u=\beta u^{*}$ where $f(u)\leq\epsilon$. Hence, GD with the Polyak step-size and $\theta_{0}=0$ ensures that when $T$ is the first iteration s.t. $f(\theta_{T})-f(u)\leq\epsilon\implies f(\theta_{T})\leq 2\epsilon$, then, for $C:=\frac{(2c-1)}{c^{2}\,\nu}=\frac{(2c-1)}{c^{2}}$, $L:=\left[\nu^{2}\,f(\theta_{0})+B\,L_{1}\,f(\theta_{0})\right]$ where $B=\frac{e^{q}-1}{q}$ and $q=L_{1}\,\left\|\theta_{0}-u\right\|$,
$f(\theta_{T})$ $\leq f(\beta u^{*})+\frac{L}{2}\left[\beta^{2}-C\,T\,\left(\frac{\epsilon}{\epsilon+f(\beta u^{*})}\right)^{2}\right]$
$f(\theta_{T})$ $\leq\epsilon+\frac{L}{2}\left[\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}-\frac{C\,T}{4}\right]$ (Since $f(\beta u^{*})\leq\epsilon$ and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$)
@@
$T\geq\frac{4\,c^{2}}{(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
In the above result, $\eta_{t}$ depends on unknown problem-dependent constants such as $u^{*}$ and $\gamma$. In the following corollary, we show that we can remove this dependence, at the expense of obtaining a worse iteration complexity.
Corollary 7.
For logistic regression on linearly separable data with margin $\gamma$ where $u^{*}$ is the corresponding max-margin solution and $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$, if $\alpha:=\frac{f(\beta u^{*})}{\epsilon}$ and if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}$, a fixed $\epsilon\in(0,f(\theta_{0}))$,, GD with the Polyak step-size ${\eta_{t}}=\frac{f(\theta_{t})-\epsilon}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ requires
$T\geq\frac{(1+\alpha)^{2}\,c^{2}}{\alpha^{2}\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
to ensure that $f(\theta_{T})\leq 2\,\epsilon$.
Proof.
The first part of the proof considers a general convex $f$ satisfies Assn. 3 with $\omega=0$. Assume $\alpha\in(0,1)$ and consider a comparator $u$ s.t. $\alpha\epsilon\leq f(u)\leq\epsilon$. We set $\eta_{t}=\frac{f(\theta_{t})-\epsilon}{c\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ and note that for all $t$, $f(\theta_{t})\geq\epsilon$.
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla f(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$ (26)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-\epsilon]$ (Using the Polyak step-size with $c>\frac{1}{2}$ to simplify the third term)
@@
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})-f(u)]$ (Since $f(u)\leq\epsilon$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon][f(\theta_{t})-f(u)]}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ (Using the Polyak step-size )
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon]^{2}}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}}$ (Since $f(\theta_{t})>\epsilon$ for all $t$. Else, we have achieved the desired sub-optimality)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{[f(\theta_{t})-\epsilon]^{2}}{c\,\nu^{2}\,[f(\theta_{t})]^{2}}$ (Using Assn. 3 with $\omega=0$)
$\implies\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-\epsilon}{f(\theta_{t})}\right)^{2}$ (Using Lemma 1 and defining $C:=\frac{(2c-1)}{c^{2}\,\nu}$)
Assuming $T$ is the first iterate s.t. $f(\theta_{T})-f(u)\leq\epsilon$, and therefore $f(\theta_{t})-f(u)\geq\epsilon$ for all $t<T$. Hence, $f(\theta_{t})\geq f(u)+\epsilon\geq(1+\alpha)\epsilon$. Therefore
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(\frac{f(\theta_{t})-\epsilon}{f(\theta_{t})}\right)^{2}$
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(1-\frac{\epsilon}{(1+\alpha)\epsilon}\right)^{2}$
$=\left\|\theta_{t}-u\right\|_{2}^{2}-C\,\left(1-\frac{1}{(1+\alpha)}\right)^{2}$
By recursing from $t=0$ to $T-1$,
$\left\|\theta_{T}-u\right\|_{2}^{2}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-CT\left(1-\frac{1}{(1+\alpha)}\right)^{2}$
$\implies f(\theta_{T})-f(u)$ $\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{2}+\frac{L}{2}\,\left[\left\|\theta_{0}-u\right\|_{2}^{2}-C\,T\left(1-\frac{1}{(1+\alpha)}\right)^{2}\right]$ (28)
Hence, for $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to guarantee that $\left\|\theta_{T}-u\right\|_{2}^{2}\leq\frac{\epsilon}{L}$. In order to guarantee this, it is sufficient to set $T$ as follows.
$T$ $\geq\frac{\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\epsilon}{L}}{C}\,\left[1+\frac{1}{\alpha}\right]^{2}$
Using the definition of $C$, we conclude that it is sufficient to set $T$ as:
$T$ $\geq\frac{c^{2}\,\nu\,\left\|\theta_{0}-u\right\|_{2}^{2}}{(2c-1)}\,\left[1+\frac{1}{\alpha}\right]^{2}$ (29)
This completes the first part of the proof. The subsequent proof is specialized to unregularized logistic regression.
$\gamma$ $:=\min_{i}\,y_{i}\langle x_{i},u^{*}\rangle$ (30)
For a scalar $\beta>0$,
For normalized data, s.t. $\left\|x_{i}\right\|\leq 1$, the logistic regression loss is convex, is uniform smooth with $L=\frac{1}{4n}\lambda_{\max}[X^{T}X]\leq 1$. Moreover, $\nu=1$. We set $\beta=\frac{1}{\gamma}\,\ln\left(\frac{1}{\epsilon}\right)$ implies that $f(\beta u^{*})\leq\epsilon$. For the $T$ defined in the theorem statement, consider two cases:
Case (II): $f(\theta_{T})>f(\beta u^{*})$. In this case, $\alpha=\frac{f(\beta u^{*})}{\epsilon}\leq 1$ and we can use the above result. In particular, we choose $u=\beta u^{*}$ and $\theta_{0}=0$ ensuring that $\left\|\theta_{0}-u\right\|=\beta$ and $f(u)\leq\epsilon$. Plugging this value in Eq. 29, we conclude that in order to ensure that $f(\theta_{T})\leq 2\,\epsilon$, it is sufficient to set $T$ as:
$T\geq\frac{(1+\alpha)^{2}\,c^{2}}{\alpha^{2}\,(2c-1)\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon}\right)\right]^{2}$
Though the above result does not require the knowledge of any problem-dependent constants, the iteration complexity scales as $O\left(\frac{1}{\alpha^{2}}\,\ln\left(\frac{1}{\epsilon}\right)^{2}\right)$, where $\alpha=\frac{f(\beta\,u^{*})}{\epsilon}$ can be arbitrarily small though non-zero.
Corollary 8.
iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
Proof.

### v3

@@ 5.2 Gradient domination with $\zeta=2$
For a fixed $\epsilon\in\left(0,\frac{\lambda_{0}}{\lambda_{1}}\right)$, if $f$ satisfies Assn. 1, 2 and 3 and Assn. 4 with $\zeta=2$, $f^{*}=0$ and if $\mu:=\min_{t\in[T]}\mu(\theta_{t})>0$, then, GD-LS with $\eta_{\max}=\infty$, requires $T\geq\frac{2}{\mu}\,\left[\lambda_{1}{f}(\theta_{0})+\lambda_{0}\,\ln\left(\frac{\lambda_{0}}{\lambda_{1}\,\epsilon}\right)\right]$ iterations to ensure that $f(\theta_{T})\leq\epsilon$.
@@ 6 SGD with Stochastic Line-search
When minimizing uniformly-smooth convex functions in the interpolation setting, Vaswani et al. (2019a) proved that SGD-SLS converges to the optimum at an $O(1/\epsilon)$ rate, matching GD and is faster than the standard $O(1/\epsilon^{2})$ rate for SGD (Bottou et al., 2018). Motivated by this and the results in Sec. 4, we analyze SGD-SLS for logistic regression which satisfies Assn. 1, 2 and 3. We first note the results in Vaswani et al. (2019a) (and subsequent papers analyzing the interpolation setting) do not directly apply to logistic regression. In particular, these results have a dependence on $\left\|\theta^{*}\right\|$ which is infinite in the logistic regression example (Orabona, 2024). Consequently, we first prove that SGD-SLS can exploit the uniform smoothness in logistic regression and prove a $O(1/\epsilon)$ rate in App. E.
Theorem 4.
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,1\right)$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ requires $T=O\left(\frac{1}{\epsilon\,\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right)$ iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq\frac{5\epsilon}{2}$.
Next, we exploit the non-uniform smoothness of logistic regression and analyze the convergence of SGD-SLS. Note that since SLS involves an Armijo line-search for one (randomly chosen) function in each iteration, we can follow the same argument as in Lemma 1 and show that the step-size in each iteration is lower-bounded i.e. ${\eta_{t}}\geq\min\left\{\eta_{\max},\frac{1}{\lambda_{0}\,+\lambda_{1}\,{f_{t}}(\theta_{t})}\right\}$. Given this result, we prove the following theorem in App. E.
Theorem 5.
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{C^{\prime}}{8}\right)$ where $C^{\prime}=O(1)$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ requires $T=O\left(\frac{n}{\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}\right)$ iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq\frac{5\epsilon}{2}$.
Note that the above $O(n\,(\ln(n/\epsilon))^{2})$ convergence rate is slower than that of GD-LS. However, since SGD-SLS has an $O(1)$ cost per iteration (as compared to the $O(n)$ cost for GD-LS), both algorithms have the same gradient complexity. In order to intuitively understand why it is difficult to prove a faster (independent of $n$) rate for SGD-SLS, first note that the stochastic setting does not allow using arbitrarily large values of $\eta_{\max}$. Second, note that individual losses $f_{i}$ in the finite sum can become much smaller than $f$. Specifically, consider iteration $t$ of SGD-SLS and consider a point $i$ such that $f_{i}(\theta_{t})\leq\delta<<\epsilon$ where $\epsilon$ is the desired sub-optimality. If this point is sampled at iteration $t$, the corresponding size of the update is $\left\|\theta_{t+1}-\theta_{t}\right\|\approx\delta\,\eta_{\max}=\nicefrac{{\delta}}{{\epsilon}}$ which can be small. On the other hand, for $i$ such that $f_{i}(\theta_{t})\geq\delta$, the size of the update is $O(1)$. Hence, if at iteration $t$, there are n1n-1 “small” losses and the algorithm samples a point uniformly at random, it has an $O(1)$ update with probability $1/n$. Hence, in expectation, the rate depends on $n$ in the worst-case.
However, note that the same SGD-SLS algorithm can achieve the rates in Thm. 4 and 5 and hence, the algorithm has an $O\left(\min\{n,\nicefrac{{1}}{{\epsilon}}\}\,(\ln(n/\epsilon))^{2}\right)$ convergence rate. We conjecture that by formalizing the above intuition, we can prove a matching lower-bound and leave this to future work. Note that the resulting gradient complexity for SGD-SLS is smaller than that of GD-LS. Interestingly, this is also smaller than the $O\left(\left(n+\nicefrac{{1}}{{\epsilon}}\right)\,\ln(1/\epsilon)\right)$ gradient complexity for variance reduced methods such as SARAH (Nguyen et al., 2017) on general uniformly-smooth convex losses.
Given the above intuition, a natural question is whether SGD-SLS can attain faster rates if the algorithm can ensure that it samples points that have a relatively large function values. We formalize this in the following theorem proved in App. E.
Theorem 6.
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{\min\left\{\frac{1}{2},C^{\prime}\right\}}{8}\right)$ where $C^{\prime}:=\frac{1}{\lambda_{1}}=\frac{1}{648}$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ and guaranteeing that for all $t\in[T]$, $\Pr\left[{f_{t}}(\theta_{t})\geq\frac{\epsilon}{2}\right]=1$ requires $T=O\left(\frac{1}{\gamma^{2}}\,\left[\ln\left(\frac{1}{\epsilon^{2}}\right)\right]^{2}\right)$ iterations to ensure that $\mathbb{E}[f(\theta_{T})]\leq\frac{3\epsilon}{2}$.
Hence, if the algorithm can ensure that the sampled function $f_{t}$} at iteration $t$ has a loss greater than $\epsilon/2$, then, SGD-SLS can indeed achieve a faster $O((\ln(1/\epsilon))^{2}$ convergence rate which is independent of $n$. Skipping updates when the sampled point has a small loss, or using a projection step to ensure that no loss becomes smaller than $\epsilon/2$ are two potential ways to ensure that $\Pr\left[{f_{t}}(\theta_{t})\geq\frac{\epsilon}{2}\right]=1$. Implementing such approaches in a computationally efficient manner and analyzing the resulting algorithm is challenging, and we leave this interesting direction to future work. Finally, we note that by combining the proof techniques in Sec. C.1 and App. E, the above convergence guarantees also hold for the stochastic Polyak step-size (Loizou et al., 2021).
@@ Acknowledgments
Corollary 5.
iterations to ensure that $f(\theta_{T})-f^{*}\leq\epsilon$.
Proof.
@@ C.1 Proofs for the Polyak Step-size
Following a proof similar to that for Thm. 2, for an arbitrary comparator $u$ s.t. $f(u)\leq f(\theta_{T})$,
@@
Theorem 8.
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for an initialization $\theta_{0}=0$, a fixed $\epsilon\in(0,f(\theta_{0}))$, GD with the Polyak step-size ${\eta_{t}}=\min\left\{\frac{f(\theta_{t})}{c\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}},\,\frac{1}{2\,\epsilon}\right\}$ for some $c>2$ requires
$T\geq\frac{\beta^{2}}{C}=\frac{64\,c^{2}}{(c-1)\,\gamma^{2}}\left[\ln\left(\frac{64\,c}{\epsilon}\right)\right]^{2}$
@@
The logistic loss on linearly separable data is convex, satisfies Assn. 1, 2 and 3 with $L_{0}=0,\,\omega=0,\,\nu=8$, and $f^{*}=0$. By Assn. 3, $\left\|\nabla f(\theta)\right\|\leq\nu\,f(\theta)$ and we can bound the Polyak step-size as:
${\eta_{t}}$ $\in\left[\min\left\{\frac{1}{c\,\nu^{2}\,f(\theta_{t})},\frac{1}{2\epsilon}\right\},\frac{1}{2\epsilon}\right].$ (23)
Using the GD update: $\theta_{t+1}=\theta_{t}-{\eta_{t}}\,\nabla f(\theta_{t})$, consider a comparator $u$ s.t. $f(u)\leq\frac{\epsilon}{\max\{c\,\nu^{2},2\}}$ and $f(u)\leq f(\theta_{t})$ for all $t\in[T]$. Assuming that $T$ is the first iteration such that $f(\theta_{T})-f(u)\leq\epsilon$, we have that,
$\left\|\theta_{t+1}-u\right\|_{2}^{2}$ $=\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,\langle\nabla f(\theta_{t}),\theta_{t}-u\rangle+{\eta_{t}}^{2}\,\left\|\nabla f(\theta_{t})\right\|_{2}^{2}$
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-2\,{\eta_{t}}\,[f(\theta_{t})-f(u)]+\frac{{\eta_{t}}}{c}\,[f(\theta_{t})]$ (Using the Polyak step-size with $c>2$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,{\eta_{t}}\,f(\theta_{t})+2\,{\eta_{t}}\,f(u)$
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\min\left\{\frac{1}{c\,\nu^{2}},\frac{f(\theta_{t})}{2\epsilon}\right\}+2\,{\eta_{t}}\,f(u)$ (Using Eq. 23)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{1}{\max\left\{c\,\nu^{2},2\right\}}+2\,{\eta_{t}}\,f(u)$ (Since $f(\theta_{t})\geq\epsilon$ for all $t<T$)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{1}{\max\left\{c\,\nu^{2},2\right\}}+\frac{f(u)}{\epsilon}$ (Since ${\eta_{t}}\leq\frac{1}{2\,\epsilon}$ from Eq. 23)
$\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\left(2-\frac{1}{c}\right)\,\frac{1}{\max\left\{c\,\nu^{2},2\right\}}+\frac{1}{\max\left\{c\,\nu^{2},2\right\}}$ (Since $f(u)\leq\frac{\epsilon}{\max\{c\,\nu^{2},2\}}$)
$=\left\|\theta_{t}-u\right\|_{2}^{2}-\underbrace{\left(1-\frac{1}{c}\right)\,\frac{1}{\max\left\{c\,\nu^{2},2\right\}}}_{:=C}$
Summing up from $t=0$ to t=T1t=T-1,
$\left\|\theta_{T}-u\right\|_{2}^{2}\leq\left\|\theta_{0}-u\right\|_{2}^{2}-CT=\left\|u\right\|_{2}^{2}-C\,Tu$ (Since $\theta_{0}=0$)
Since $f$ is $1$ uniformly smooth, using Lemma 4 with $M=f(\theta_{0})$, we get
$f(\theta_{T})-f(u)\leq\frac{\epsilon}{2}+\underbrace{[\nu^{2}\,f(\theta_{0})+1]}_{:=L}\frac{\left\|\theta_{T}-u\right\|_{2}^{2}}{2}\leq\frac{\epsilon}{2}+L\,\left[\left\|u\right\|_{2}^{2}-C\,T\right]$
To ensure that $f(\theta_{T})-f(u)\leq\epsilon$, it is sufficient to set
$T\geq\frac{\left\|u\right\|_{2}^{2}}{C}.$ (24)
In order to bound $\left\|u\right\|$, we define $u^{*}$ to be the max-margin solution i.e. $\left\|u^{*}\right\|=1$ and $\gamma$ to be the corresponding margin, i.e.
@@
Consider $u=\beta\,u^{*}$, for a scalar $\beta=\frac{1}{\gamma}\,\ln\left(\frac{\max\{c\nu^{2},2\}}{\epsilon}\right)$,
$f(u)$ $=\frac{1}{n}\,\sum_{i=1}^{n}\ln(1+\exp(-y_{i}\langle x_{i},\beta u^{*}\rangle))\leq\frac{1}{n}\,\sum_{i=1}^{n}\exp(-y_{i}\langle x_{i},u^{*}\rangle)\leq\exp(-\beta\gamma)=\frac{\epsilon}{\max\{c\nu^{2},2\}}$ (26)
This satisfies the requirement on $f(u)$. For logistic regression, we $\nu=8$ and since $c>2$, therefore $\max\{c\,\nu^{2},2\}=c\nu^{2}$. Using this to bound $T$, we get that,
$T$ $\geq\frac{\beta^{2}}{C}=\frac{c^{2}\,\nu^{2}}{(c-1)\,\gamma^{2}}\left[\ln\left(\frac{c\,\nu^{2}}{\epsilon}\right)\right]^{2}$
@@
Finally, we conclude that after $T=\frac{\beta^{2}}{C}=\frac{c^{2}\,\nu^{2}}{(c-1)\,\gamma^{2}}\left[\ln\left(\frac{c\,\nu^{2}}{\epsilon}\right)\right]^{2}$ iterations we have $f(\theta_{T})-f(u)\leq\epsilon$ and since $f(u)\leq\epsilon$ then $f(\theta_{T})\leq 2\,\epsilon$